OpenStudy (kpop_fan):

help please!! A box has 1 red marble, 1 blue marble, and 4 green marbles of the same shape and size. Dan draws a marble randomly from the box, replaces it, and then draws another marble randomly. What is the probability of drawing two green marbles in a row? Explain your answer

OpenStudy (witty_roger):

(4/6) * (4/6) = 4/9

OpenStudy (crabbyoldgamer):

Probability is desired/total: on the first draw that's 4/6: 4 green, 6 total. Since the marble is replaced, the probability on the second draw is also 4/6. The probability of green then green is 4/6 * 4/6 = 16/36 = 4/9